OpenStudy (anonymous):

Twelve is the geometric mean of 9 and what number? 16 3 24 4

OpenStudy (anonymous):

\(\sqrt{3\times3\times4\times4}=12 \) 16

OpenStudy (anonymous):

In symbols, your claim is as follows.\[\sqrt{9\cdot x}=12\]